1/2

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

The Werewolf  Varulven  El Hechicero Der Werwolf

translated by Alexander Gross translated by Sven Collberg translated by Hans Eiseneck

A Werewolf, troubled by his

name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

En varulv lopp från

fru m.m. en vacker afton samt

begav sig till en bymagi-

sters grav och bad: “Var snäll

och konjugera.

Un hechicero, con

esposa vino un día a la fo-

sa de su maestro, y al

profesor pidió: “¿Me cuen-

tas? ¡Por favor!”

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschulleh-

rers Grab und bat ihn: �Bitte, beuge mich!“

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready. mig!” Byskolläraren

stod upp på bleckna-

mnsskyltens mässingknopp

med svar till den,

som satt på pass med tåligt korslagd

tass på tass.

El profesor sin va-

cilar subió y comenzó a

hablar. Desde allá de su le-

trero se dirigió al hechi-

cero:

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messingknauf und sprach zum Wolf, der seine Pfoten geduldig kreuzte vor dem Toten:

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singular-

ly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

“Jag var-ulv 1:a pers.

i sing., du, han, hon, den,

det var-ulv. Så vi voro-ulv, I voren.

På de voro-ulv kom in-

genting.”

“El hechiuno” le

explica, “el hechidos” de-

spués indica, “el hechitres” dice

al fin, “con eso basta ¡chi-

quitín!”

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Genitiv

sodann, dem Werwolf, Dativ, wie man’s nennt, den Werwolf, - damit hat’s ein End!“

“I know that–I’m no mental

cripple– The future form and parti-

ciple Are what I crave,” the beast

replied. The scholar paused–again

he tried:

Varulven myste glad

åt hur pers. hade följts av

sing. och plur. “Fast”, bad han, “kan

det inte ges ens en enda pers. i tem-

pus presens?”

El hechicero, con-

movido por lo que capta su

oído, al maestro ruega,

encantado, que le adjunte un

quebrado.

Dem Werwolf schmeichelten die Fälle er rollte seine Au-

genbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl noch!“

…  …  …  …

2/2

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-

wolf?’ Utterly wrong! Such words are wounds

beyond all suture– I’m sorry, but you have no

future.”

Bekänna måste vår

magister, att därvidlag fanns

vissa brister. En varulv sine imper-

fecto var contradictio in

adjecto.

Pero el profesor

confiesa que nada sabe con

certeza: “Hay muchos he-

chos. Hechiceros

únicamente hay enteros.”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

The Werewolf knew better–

his sons still slept At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

Varulven satt med

tårad lins. Han visste att han

fanns och finns. Därtill kom också fru

m.m. som även måste exi-

stera.

El brujo se volvió

muy triste: “Mi cara mitad,

¿no la viste?” Ya que no era sa-

bio con gratitud se de-

spidió.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Swedish by Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into hispana by Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)